OpenStudy (anonymous):

MEDALS!!!! PLEASE HELP....A jar contains 5 red jelly beans, 9 yellow jelly beans, and 16 orange jelly beans. Suppose that each jelly bean has an equal chance of being picked from the jar. If a jelly bean is selected at random from the jar, what is the probability that it is NOT red? 1/5, 2/3, 4/5 or 5/6.....

OpenStudy (anonymous):

The answer is 5/6

OpenStudy (anonymous):

thank you. I am not good a probability at all.

OpenStudy (anonymous):

Since you are not including red jelly probability that it is NOT red = (no. of yellow +no. orange)/total number of jellies = (9+16)/(5+9+16)= 25/30=5/6

OpenStudy (anonymous):

@JANELLSM98

OpenStudy (anonymous):

that makes more sense. seeing it written out.